# Dependency-aware Maximum Likelihood Estimation for Active Learning

### v1

@@ 6 Conclusion and Future Works
In this paper, we introduced Dependency-aware Maximum Likelihood Estimation (DMLE) for the active learning framework, a novel approach that highlights the incompatibility of the i.i.d. data assumption in conventional parameter estimation methods with the active learning paradigm. Specifically, DMLE preserves the sample dependency that would otherwise be overlooked in conventional independent MLE (IMLE), ensuring they are properly taken into account in the estimation. We examined the impact of this dependency term across various acquisition functions, sample selection strategies, and sample sizes, evaluating its effect on multiple benchmark datasets with varying model complexities. Our experiments revealed a significant improvement in average test accuracy when using DMLE compared to IMLE, underlining the critical role of accounting for sample dependencies in model parameter estimation. Future work directions include exploring new modeling strategies to approximate the conditional distribution of sample dependencies.
@@ Appendix A Stochastic Batch Selection
    $P(\mathbf{s}_{t+1}\mid D_{t};\theta)=\frac{1}{Z_{t}}\prod_{i=1}^{k}e^{\beta a(x_{i},D_{t},\theta)}~{}\text{ where }~{}Z_{t,i}\equiv\sum_{x\in U_{t}\setminus\{x_{1},\ldots,x_{i-1}\}}e^{\beta a(x,D_{t},\theta)}.$
@@
    $P(\mathbf{s}_{t+1}\mid D_{t};\theta)=\frac{1}{Z_{t}}\prod_{i=1}^{k}a(x_{i},D_{t},\theta)^{\beta}~{}\text{ where }~{}Z_{t,i}\equiv\sum_{x\in U_{t}\setminus\{x_{1},\ldots,x_{i-1}\}}a(x,D_{t},\theta)^{\beta}.$
@@
    $P(\mathbf{s}_{t+1}\mid D_{t};\theta)=\frac{1}{Z_{t}}\prod_{i=1}^{k}r_{i}^{-\beta}~{}\text{ where }~{}Z_{t,i}\equiv\sum_{x\in U_{t}\setminus\{x_{1},\ldots,x_{i-1}\}}r^{-\beta}$
@@ Appendix C Proofs
The expected log-likelihood of the learned model $\ \hat{\theta}$ over the true data distribution $P^{true}$ is given by:
$\mathbb{E}_{(x^{\prime},y^{\prime})\sim P^{true}}\Bigl{[}\ln L(\hat{\theta})\Bigr{]}$ $=\mathbb{E}_{(x^{\prime},y^{\prime})\sim P^{true}}\Bigl{[}\ln P(y^{\prime}\mid x^{\prime};\hat{\theta})+\ln P(x^{\prime}\mid D_{t};\hat{\theta})\Bigr{]}.$ (36)
@@
$\sum_{x^{\prime}}P(x^{\prime})\sum_{y^{\prime}}P(y^{\prime}\mid x^{\prime})\Bigl{[}\ln P(y^{\prime}\mid x^{\prime};\hat{\theta})+\ln P(x^{\prime}\mid D_{t};\hat{\theta})\Bigr{]}.$ (37)
@@
$\sum_{x^{\prime}}P(x^{\prime})\sum_{y^{\prime}}P(y^{\prime}\mid x^{\prime})\ln P(y^{\prime}\mid x^{\prime};\hat{\theta})+\sum_{x^{\prime}}P(x^{\prime})\ln P(x^{\prime}\mid D_{t};\hat{\theta}).$ (38)
@@
$\mathbb{E}_{(x^{\prime},y^{\prime})\sim P^{true}}\Bigl{[}\ln L(\hat{\theta})\Bigr{]}=-D_{\text{KL}}(P(y^{\prime}\mid x^{\prime})\parallel P(y^{\prime}\mid x^{\prime};\hat{\theta}))-D_{\text{KL}}(P(x^{\prime})\parallel P(x^{\prime}\mid D_{t};\hat{\theta})).$ (40)
@@
Appendix E Additional Results
To further analyze the importance of taking the sample dependencies into account during model parameter estimation in active learning, we plot the test accuracy plots with ±$\pm 1$ standard deviation for four more datasets, e.g., Mnist, Fashion-Mnist (FMnist), Emnist, and Cifar-10, where for each we utilize four different sample selection strategies explained in Section 3.2. In Figure 7, we observe the performance improvement with DMLE when compared with IMLE in most cases. For the cases where DMLE doesn’t outperform IMLE, we believe further hyperparameter tuning for $\beta$ might lead to better performance. For all experiments here, we use $\beta=1$ as suggested by Kirsch et al. (2021).
MNIST-SSMS
MNIST-SPS
MNIST-SSRS
MNIST-Top-$k$
CIFAR10-SSMS
CIFAR10-SPS
CIFAR10-SSRS
CIFAR10-Top-k
(i) FMNIST-SSMS
(j) FMNIST-SPS
(k) FMNIST-SSRS
(l) FMNIST-Top-$k$
(m) EMNIST-SSMS
(n) EMNIST-SPS
(o) EMNIST-SSRS
(p) EMNIST-Top-$k$
Figure 7: The average test accuracy comparison with ±$\pm 1$ standard deviation for DMLE and IMLE over cycles for Mnist, Cifar-10, FMnist, and Emnist datasets for different sample selection strategies, namely Stochastic Softmax Sampling(SSMS), Stochastic Power Sampling(SPS), Stochastic Soft-rank Sampling(SSRS), and Top-$k$ Sampling where sample selection size is $k=1$. In the earlier cycles, DMLE was observed to outperform IMLE in all cases except for FMnist with SSRS and Top-k which is highly advantageous in active learning.
IRIS-SSMS (k=1)
IRIS-SPS (k=1)
IRIS-SSRS (k=1)
IRIS-Top-$k$ (k=1)
IRIS-SSMS (k=5)
IRIS-SPS (k=5)
IRIS-SSRS (k=5)
IRIS-Topk-$k$ (k=5)
IRIS-SSMS (k=10)
IRIS-SPS (k=10)
IRIS-SSRS (k=10)
IRIS-Top-$k$ (k=10)
Figure 8: The average test accuracy comparison with ±$\pm 1$ standard deviation for DMLE and IMLE over cycles for Iris for different sample selection strategies, namely Stochastic Softmax Sampling(SSMS), Stochastic Power Sampling(SPS), Stochastic Soft-rank Sampling(SSRS), and Top-$k$ Sampling where sample selection sizes are $k=1$, $k=5$, and $k=10$. As the sample selection size increases, the significance of using DMLE over IMLE becomes more acute.
Table 3: Comparison of mean classification test accuracies ±$\pm 1$ standard deviation for different sample selection sizes ($k$) during cycles and different sample selection strategies where number of samples in the labeled set $N_{L}=100$. We use BALD for the acquisition function and SSMS (Stochastic Softmax Sampling), SPS (Stochastic Power Sampling), SSRS (Stochastic Soft-rank Sampling), and Top-$k$ sampling for the sample selection strategy. The bold text highlights the higher mean accuracy and lower standard deviation when comparing DMLE and IMLE under various sampling schemes.
Dataset $k$ SSMS SPS SSRS Top-$k$
DMLE IMLE DMLE IMLE DMLE IMLE DMLE IMLE
MNIST 1 $\mathbf{0.77\pm 0.02}$ $0.77\pm 0.03$ $0.77\pm 0.03$ $\mathbf{0.78\pm 0.03}$ $\mathbf{0.82\pm 0.02}$ $0.80\pm 0.03$ $0.80\pm 0.05$ $\mathbf{0.80\pm 0.04}$
10 $\mathbf{0.48\pm 0.04}$ $0.47\pm 0.04$ $\mathbf{0.48\pm 0.02}$ $0.47\pm 0.01$ $\mathbf{0.46\pm 0.01}$ $0.45\pm 0.03$ $\mathbf{0.44\pm 0.01}$ $0.42\pm 0.04$
Table 4: Comparison of mean classification test accuracies ±$\pm 1$ standard deviation for different sample selection sizes ($k$) during cycles and different sample selection strategies where number of samples in the labeled set $N_{L}=100$. We use Least Confidence for the acquisition function and SSMS (Stochastic Softmax Sampling), SPS (Stochastic Power Sampling), SSRS (Stochastic Soft-rank Sampling), and Top-$k$ sampling for the sample selection strategy. The bold text highlights the higher mean accuracy and lower standard deviation when comparing DMLE and IMLE under various sampling schemes.
Dataset $k$ SSMS SPS SSRS Top-$k$
DMLE IMLE DMLE IMLE DMLE IMLE DMLE IMLE
5 $\mathbf{0.79\pm 0.03}$ $0.77\pm 0.02$ $\mathbf{0.77\pm 0.04}$ $0.76\pm 0.03$ $\mathbf{0.77\pm 0.02}$ $0.74\pm 0.04$ $\mathbf{0.68\pm 0.06}$ $0.66\pm 0.05$
10 $\mathbf{0.77\pm 0.03}$ $\mathbf{0.76\pm 0.03}$ $0.76\pm 0.02$ $\mathbf{0.77\pm 0.03}$ $\mathbf{0.75\pm 0.04}$ $0.73\pm 0.05$ $\mathbf{0.62\pm 0.08}$ $0.54\pm 0.07$
EMNIST 1 $\mathbf{0.35\pm 0.02}$ $0.35\pm 0.03$ $0.37\pm 0.03$ $\mathbf{0.37\pm 0.01}$ $\mathbf{0.37\pm 0.02}$ $0.35\pm 0.04$ $0.31\pm 0.02$ $\mathbf{0.33\pm 0.02}$
10 $\mathbf{0.51\pm 0.03}$ $0.47\pm 0.04$ $\mathbf{0.50\pm 0.02}$ $0.48\pm 0.02$ $0.44\pm 0.07$ $\mathbf{0.46\pm 0.04}$ $\mathbf{0.33\pm 0.10}$ $0.30\pm 0.09$
5 $\mathbf{0.25\pm 0.02}$ $0.23\pm 0.04$ $\mathbf{0.24\pm 0.03}$ $0.23\pm 0.03$ $\mathbf{0.22\pm 0.04}$ $\mathbf{0.22\pm 0.04}$ $\mathbf{0.20\pm 0.02}$ $0.19\pm 0.04$
10 $0.23\pm 0.03$ $\mathbf{0.23\pm 0.02}$ $\mathbf{0.23\pm 0.03}$ $\mathbf{0.23\pm 0.03}$ $\mathbf{0.21\pm 0.04}$ $0.20\pm 0.02$ $\mathbf{0.20\pm 0.02}$ $\mathbf{0.20\pm 0.02}$
CIFAR-10 1 $\mathbf{0.23\pm 0.02}$ $0.23\pm 0.03$ $\mathbf{0.24\pm 0.03}$ $0.23\pm 0.02$ $\mathbf{0.22\pm 0.02}$ $0.21\pm 0.02$ $\mathbf{0.20\pm 0.05}$ $0.20\pm 0.03$
5 $\mathbf{0.22\pm 0.04}$ $0.18\pm 0.02$ $\mathbf{0.23\pm 0.04}$ $0.20\pm 0.03$ $\mathbf{0.20\pm 0.02}$ $0.19\pm 0.03$ $\mathbf{0.19\pm 0.02}$ $0.18\pm 0.02$
Appendix F Experimental Timing Results

### v2

@@ 6 Conclusion and Future Works
In this paper, we introduced Dependency-aware Maximum Likelihood Estimation (DMLE) for the active learning framework, a novel approach that highlights the incompatibility of the i.i.d. data assumption in conventional parameter estimation methods with the active learning paradigm. Specifically, DMLE preserves the sample dependency that would otherwise be overlooked in conventional independent MLE (IMLE), ensuring they are properly taken into account in the estimation. We examined the impact of this dependency term across various acquisition functions, sample selection strategies, and sample sizes, evaluating its effect on multiple benchmark datasets with varying model complexities. Our experiments revealed a significant improvement in average test accuracy when using DMLE compared to IMLE, underlining the critical role of accounting for sample dependencies in model parameter estimation.
A promising direction for future work is the theoretical characterization of the dependency term’s distribution under Top-$k$ selection strategies, beyond the current empirical analysis we provide. This would lead to deeper insights into the approximation behavior of DMLE and help identify scenarios where its benefits may diminish. Additionally, examining DMLE’s performance across varying uncertainty and diversity conditions represents a valuable line of research, with potential to inform improvements in the acquisition process. Another promising direction is exploring connections between DMLE and submodular optimization frameworks for large-batch selection. Prior work has shown that submodularity can effectively guide the selection of informative and diverse subsets under efficiency constraints (Golovin & Krause, 2017). Adapting DMLE within such a combinatorial framework may help scale its benefits to larger selection sizes while maintaining tractable optimization.
@@ Appendix A Stochastic Batch Selection
[…]
@@
[…]
@@
[…]
@@ Appendix C Proofs
[…]
@@
[…]
@@
[…]
@@
$).\displaystyle\mathbb{E}_{(x^{\prime},y^{\prime})\sim P^{true}}\Bigl[\ln L(\hat{\theta})\Bigr]=-D_{\text{KL}}(P(y^{\prime}\mid x^{\prime})\parallel P(y^{\prime}\mid x^{\prime};\hat{\theta}))-D_{\text{KL}}(P(x^{\prime})\parallel P(x^{\prime}\mid D_{t};\hat{\theta}$)). (40)
@@
Appendix E Experimental Timing Results
@@
As the difference between DMLE and IMLE lies in the objective function, we also plot the loss processing times per cycle to better observe the computational expense introduced by the dependency term in DMLE, as shown in Figure 7. This experiment, conducted on the MNIST dataset collecting 500 samples with sample selection size $k=1$, shows that the loss processing time increases for both methods across cycles as more samples are labeled. Although the difference between DMLE and IMLE grows over time, it remains relatively small in practice.
Figure 7: Per-cycle loss processing time for DMLE and IMLE during active learning on MNIST (500 cycles, sample selection size $k=1$). This plot illustrates how the computational cost evolves across cycles, emphasizing the impact of DMLE’s dependency-aware loss formulation. While both methods exhibit increased processing times as the labeled set grows, DMLE shows a steeper rise due to the additional overhead of modeling dependencies. Nonetheless, within the context of active learning, where the objective is to achieve high performance with a limited number of labeled samples, the overall runtime remains practical, demonstrating that the additional computational cost of DMLE is manageable in typical active learning scenarios.
Appendix F Additional Results
To further analyze the importance of taking the sample dependencies into account during model parameter estimation in active learning, we plot the test accuracy plots with $\pm 1$ standard deviation for four more datasets, e.g., Mnist, Fashion-Mnist (FMnist), Emnist, and Cifar-10, where for each we utilize four different sample selection strategies explained in Section 3.2. In Figure 8, we observe the performance improvement with DMLE when compared with IMLE in most cases. For the cases where DMLE doesn’t outperform IMLE, we believe further hyperparameter tuning for $\beta$ might lead to better performance. For all experiments here, we use $\beta=1$ as suggested by Kirsch et al. (2021).
MNIST-SSMS
MNIST-SPS
MNIST-SSRS
MNIST-Top-$k$
CIFAR10-SSMS
CIFAR10-SPS
CIFAR10-SSRS
CIFAR10-Top-k
(i) FMNIST-SSMS
(j) FMNIST-SPS
(k) FMNIST-SSRS
(l) FMNIST-Top-$k$
(m) EMNIST-SSMS
(n) EMNIST-SPS
(o) EMNIST-SSRS
(p) EMNIST-Top-$k$
Figure 8: The average test accuracy comparison with $\pm 1$ standard deviation for DMLE and IMLE over cycles for Mnist, Cifar-10, FMnist, and Emnist datasets for different sample selection strategies, namely Stochastic Softmax Sampling(SSMS), Stochastic Power Sampling(SPS), Stochastic Soft-rank Sampling(SSRS), and Top-$k$ Sampling where sample selection size is $k=1$. In the earlier cycles, DMLE was observed to outperform IMLE in all cases except for FMnist with SSRS and Top-k which is highly advantageous in active learning.
IRIS-SSMS (k=1)
IRIS-SPS (k=1)
IRIS-SSRS (k=1)
IRIS-Top-$k$ (k=1)
IRIS-SSMS (k=5)
IRIS-SPS (k=5)
IRIS-SSRS (k=5)
IRIS-Topk-$k$ (k=5)
IRIS-SSMS (k=10)
IRIS-SPS (k=10)
IRIS-SSRS (k=10)
IRIS-Top-$k$ (k=10)
Figure 9: The average test accuracy comparison with $\pm 1$ standard deviation for DMLE and IMLE over cycles for Iris for different sample selection strategies, namely Stochastic Softmax Sampling(SSMS), Stochastic Power Sampling(SPS), Stochastic Soft-rank Sampling(SSRS), and Top-$k$ Sampling where sample selection sizes are $k=1$, $k=5$, and $k=10$. As the sample selection size increases, the significance of using DMLE over IMLE becomes more acute.
Table 4: Comparison of mean classification test accuracies $\pm 1$ standard deviation for different sample selection sizes ($k$) during cycles and different sample selection strategies where number of samples in the labeled set $N_{L}=100$. We use BALD for the acquisition function and SSMS (Stochastic Softmax Sampling), SPS (Stochastic Power Sampling), SSRS (Stochastic Soft-rank Sampling), and Top-$k$ sampling for the sample selection strategy. The bold text highlights the higher mean accuracy and lower standard deviation when comparing DMLE and IMLE under various sampling schemes.
Dataset $k$ SSMS SPS SSRS Top-$k$
DMLE IMLE DMLE IMLE DMLE IMLE DMLE IMLE
MNIST 1 $\mathbf{0.77\pm 0.02}$ $7\pm 0.03 0.77±0$$7\pm 0.03$ $\mathbf{0.78\pm 0.03}$ $\mathbf{0.82\pm 0.02}$ $0.80\pm 0.03$ $0.80\pm 0.05$ $\mathbf{0.80\pm 0.04}$
[…]
Table 5: Comparison of mean classification test accuracies $\pm 1$ standard deviation for different sample selection sizes ($k$) during cycles and different sample selection strategies where number of samples in the labeled set $N_{L}=100$. We use Least Confidence for the acquisition function and SSMS (Stochastic Softmax Sampling), SPS (Stochastic Power Sampling), SSRS (Stochastic Soft-rank Sampling), and Top-$k$ sampling for the sample selection strategy. The bold text highlights the higher mean accuracy and lower standard deviation when comparing DMLE and IMLE under various sampling schemes.
Dataset $k$ SSMS SPS SSRS Top-$k$
DMLE IMLE DMLE IMLE DMLE IMLE DMLE IMLE
[…]
10 $\mathbf{0.51\pm 0.03}$ $0.47\pm 0.04$ $\mathbf{0.50\pm 0.02}$ $8\pm 0.02 0.44±0$$4\pm 0.07$ $\mathbf{0.46\pm 0.04}$ $\mathbf{0.33\pm 0.10}$ $0.30\pm 0.09$
[…]
10 $0.23\pm 0.03$ $\mathbf{0.23\pm 0.02}$ $\mathbf{0.23\pm 0.03}$ $\mathbf{0.23\pm 0.03}$ $\mathbf{0.21\pm 0.04}$ $\pm 0.02 0.20±0.02\mathbf$$\pm 0.02$} $\mathbf{0.20\pm 0.02}$
[…]
5 $\mathbf{0.22\pm 0.04}$ $0.18\pm 0.02$ $\mathbf{0.23\pm 0.04}$ $\pm 0.03 0.20±0.02\mathbf$$\pm 0.02$} $0.19\pm 0.03$ $\mathbf{0.19\pm 0.02}$ $0.18\pm 0.02$
@@ Appendix H Dataset/Model Licenses and Sources
Appendix I Ablation Study for Dependency Approximation under Top-$k$ Sampling
The primary motivation for choosing the Stochastic Softmax approximation is its computational simplicity. As shown in equation 14, this approximation enables the reuse of already computed acquisition scores without requiring additional operations. In contrast, the Stochastic Power method (equation 16) involves computing the logarithm of the scores, while the Stochastic Soft-rank method (equation 18) requires both sorting and logarithmic computation. These additional steps introduce extra computational overhead during query selection.
To assess the practical impact of this choice, we conducted empirical comparisons using these distributions to model the dependency under Top-$k$ selection. In Figure 10, we present results on the MNIST, Iris, Reuters, and Tiny-ImageNet datasets, which were selected to reflect varying data modalities, complexities, and sample selection sizes. The results show that the Stochastic Softmax consistently outperforms the other approximations in terms of average test accuracy under Top-$k$ sampling.
This empirical evidence supports the choice of a softmax-based approximation, highlighting its efficiency and effectiveness.
MNIST
IRIS
REUTERS
TINY IMAGENET
Figure 10: Comparison of average test accuracy (±1 standard deviation) using different distribution approximations for the dependency term with Top-$k$ selection: Stochastic Softmax (SSMS), Stochastic Power (SPS), and Stochastic Soft-rank (SSRS). These methods were evaluated during model parameter estimation on MNIST, Iris, Reuters, and Tiny ImageNet datasets. The results demonstrate the effect of each approximation across diverse data modalities and complexities, with sample selection sizes of $k=1$ for the first three datasets and $k=5$ for Tiny ImageNet.
Appendix J Glossary
This section presents a glossary of variables and notations used throughout the paper, provided in Table 6.
Term / Symbol Definition
$\mathcal{X}\subset\mathbb{R}^{d_{x}}$ Input feature space of dimension $d_{x}$.
$\mathcal{Y}=\{1,\ldots,K\}$ Label space with $K$ possible classes.
$x$ Input sample from the feature space $\mathcal{X}$.
$y$ Corresponding label of sample $x$, belonging to $\mathcal{Y}$.
$U_{t}=\{x_{i}\}_{i=1}^{N_{U}^{t}}$ Unlabeled dataset at cycle $t$, containing $N_{U}^{t}$ samples.
$D_{t}=\{(x_{i},y_{i})\}_{i=1}^{N_{L}^{t}}$ Labeled dataset at cycle $t$, containing $N_{L}^{t}$ labeled samples.
$N_{U}^{t}$ Number of unlabeled samples at cycle $t$.
$N_{L}^{t}$ Number of labeled samples at cycle $t$.
$S_{t+1}$ Batch of samples selected for labeling at cycle $t$.
$\theta\in\mathbb{R}^{d_{\theta}}$ Model parameters of dimension $d_{\theta}$.
$a(x,D,\theta)$ Acquisition function assigning a score to an unlabeled sample $x$ based on labeled data $D$, model parameters $\theta$.
$f_{\theta}:\mathcal{X}\to\{0,1\}^{K}$ Prediction function (e.g., deep neural network) parameterized by $\theta$, outputs scores for $K$ classes.
$t$ Current cycle.
Table 6: Glossary of terms and symbols used in the active learning framework.